# AI-coupled HPC Workflow Applications, Middleware and Performance

### v1

@@ 4.3.2 Multistage pipeline
In addition to dataflow across different “executables”, this gives rise to numerous performance issues. For example, the concurrent execution of multiple different stages with different computational costs and resource requirements results in a significant load balancing and workload management challenge. Furthermore, suppose AI models are used in conjunction with or as a substitute for each stage. In that case, many of the problems intrinsic to other motifs arise (e.g., dataflow performance bottlenecks in the Steering motif).
@@ 4.3.3 Inverse design
Inverse design is a technique of calculating the causal factors that produced them from a set of observations. Such approaches in various domains such as the search for catalysts [110], nanophotonics [111], and material design [112]. Performance in this motif is primarily influenced by how representative the model can capture the causal factors that produced it. This is prone to vary based on the model architecture, computational cost, and ability to work with high data rates. High-energy X-ray Diffraction Microscopy (HEDM) relies on the high-precision (sub-pixel) position of Bragg diffraction peaks. The computational cost of conventional methods becomes a bottleneck, especially for in-situ experiments with complex detector technologies. BraggNN [113] is a neural network designed to localize Bragg peak positions more efficiently than a traditional conventional pseudo-Voigt peak fitting method in extracting the peak information. A key challenge is enabling trained models to accommodate high data rates from emerging synchrotron facilities  [114].
@@ 4.3.4 Digital replica
The Digital Replica motif describes the integration of simulations with machine learning models, either in cognitive simulations (CogSim) or digital twins. The performance depends primarily on how the model is integrated into the simulation (coupling type) and compute heterogeneity. As mentioned earlier, this falls into two approaches: (1) in-memory or (2) distributed (via inference server). Jia et al. [115] used an in-memory inference approach via the TensorFlow C API to integrate a molecular dynamics (MD) simulation with a machine-learned model. Their model can efficiently scale up to the entire Summit supercomputer, achieving 91 PFLOPS in double precision and 162/275 PFLOPS in mixed-single/half-precision. For this reason, they won the Gordon Bell Prize in 2020.
@@
4.3.5 Distributed models and dynamic data
@@ 4.3.6 Adaptive execution for training large AI models
We discuss this motif for three types of implementations: Hyperparameter Optimization (HPO), Neural Architecture Search (NAS), and training LLMs. HPO is typically used to optimize a set of training hyperparameters, such as batch size, learning rate, etc.; this is typically performed by training several neural networks in an asynchronous parallel fashion. Optimization methods generally fall into either random search or Bayesian camps, with novel augmentations such as Hyperband designed to speed up the evaluations by adaptive resource allocation and early stopping [119]. Tools such as KerasTuner may be run distributed using a chief-worker strategy to scale such methods across many nodes on HPC [120].
@@
LLMs are particularly challenging to train primarily because the models are generally much larger than what can fit within the memory of a single GPU, and the models need to be trained on colossal amounts of data. For example, GPT-3 has 175 billion trainable parameters and is trained on 499 billion tokens [126]; it has been estimated that it would take 288 years to train GPT-3 on a single V100 GPU [127]. Considerable developments in data and model parallelism have enabled significant speedup by using data-based parallelism methods, such as Horovod, and a combination of pipeline and tensor model parallelism [128]. Using such a triad strategy of data, tensor, and pipeline model parallelism along with DeepSpeed by Microsoft, Narayan et al. [127] have demonstrated the ability to train an LLM with one trillion parameters at 502 PF/s on 3,072 GPUs. Additionally, with the increasing number of non-traditional AI accelerators, training LLMs require approaches to leverage the hardware. Emani et al. [129] provided a comparative evaluation of training LLM models on diverse AI accelerators. Recent studies have also further explored efficient training strategies and performance implications of up to a trillion-parameter LLMs on leadership-class supercomputers [130, 131]. Achieving such a feat required overcoming GPU memory limitations and minimizing inter-node communication latencies using a combination of pipeline and tensor model parallelism.
@@ 5 Open challenges and opportunities
In this section, we identify two main challenges in deploying AI-coupled HPC workflows at scale: (i) the need for co-designing facilities, middleware, and benchmarks that address the performance of both traditional HPC simulations and emerging AI algorithms and (ii) the need to design efficient interfaces that would allow users to describe their workflow needs and their AI-HPC interaction. In light of these challenges, this section aims to reveal open opportunities in each area for the future advancement of AI-HPC workflows.
@@ 5.2 Data and I/O Challenge
HPC platforms have settled on POSIX-compliant parallel file systems as the primary storage engine and interface. Even for an object-based solution such as DAOS, a POSIX shim layer is still needed to meet the requirement of more traditional modes in scientific applications. However, large, sequential read-and-write is no longer the dominant workload pattern for many data-intensive AI-driven workflows. Instead, small, random read/write is much more prevalent – which traditional POSIX-based parallel file systems have struggled to keep pace with. The nominal approach is to provision a node-local storage layer, where a small amount of SSD-backed devices are positioned close to the compute blades and generally outside the storage network used by the parallel file system. The challenge is that decoupling from primary storage necessitates data staging, which can be achieved through policy-driven data tiering or workflow support to streamline the process. The current solution space is not mature enough and thus renders this a particular bottleneck to overcome for data-intensive workflow integration. Projects such as NoPFS and SmartRedis discussed in Section 3 are attempts in the right direction. More recent OLCF-6 RFP [144] calls for an AI-optimized storage system and highlights the challenges and needs in this domain.
@@ 5.3 Benchmarking
To effectively evaluate the performance of AI-HPC workflows, benchmarking them on diverse use case scenarios/motifs is essential. Such benchmarks could reveal potential performance bottlenecks and, hence, lead to optimized solutions. Several efforts in this direction include MLPerf [145], HPCAI500 [146], SAIBench from BenchCouncil [147], and HPL-AI [132]. MLPerf [145], developed by the MLCommons consortium of industry, academia, and research laboratories, is a state-of-the-art benchmark suite that includes various machine learning tasks. The MLPerf HPC benchmark [148] has been developed to represent scientific machine learning workloads typical of traditional HPC environments. There is also an ongoing effort to include AI-HPC workflow benchmarks in this suite in collaboration with the MLCommons Science group [149].
@@ 5.4 Cross-cutting: Interoperability, Energy-Efficiency and Resource Allocation
AI-coupled HPC workflows often require a high-level of interoperability, typically entailing connecting/interfacing with multiple endpoints for data, AI models, and workflow integration. There are two emerging patterns of choice: (1) REST API [151] and (2) message queue-based approach [152]. Each has pros and cons; currently, no widely accepted interfacing mechanisms exist. Further, time-sensitive patterns and long-term campaign patterns, as highlighted by the IRI blueprint report [73], demonstrate the need for interfacing beyond a single institution’s boundary, which brings a host of related challenges such as federated IDs, authentication, authorization, access control, etc., to implement a true end-to-end workflow integration. These challenges are more pronounced in motifs connecting experiments with HPC simulations (e.g., Distributed Models and Inverse Design) and in supporting Steering and Multistage Pipeline motifs, which is only partially addressed by specialized frameworks such as MuMMI or DeepDriveMD, and more general frameworks such as SmartSim (ref. Section 3).
The importance of energy efficiency in high-performance computing (HPC) has intensified due to the hard limits on maximum power capacity. A critical observation in most AI-HPC workflows is their relatively low utilization of peak performance, which may be due to data movement or other I/O activities [94]. This sub-optimal utilization prompts a vital question: what should be the ultimate goal in this context? The answer might lie in optimizing how one could optimize productivity (Science/Watt) rather than raw performance (FLOP/Watt). This optimization raises a further dilemma: should the focus be on time-to-solution or throughput? One approach could be accelerating processes through high concurrency levels and GPU utilization. Alternatively, engaging GPUs less frequently might be more power-efficient, leading to a trade-off of lower throughput for energy conservation, albeit with longer processing times [153]. Other research areas for improving the energy efficiency of AI-coupled HPC workflows involve dynamic scheduling policies, data locality, ensemble pruning of unnecessary runs, and optimizations of convolution implementations [154]. Benchmarking AI-coupled HPC workflows will be key for understanding how to optimize energy efficiency [155].
@@ 6 Summary and Discussion
This paper surveyed the extensive landscape of AI-driven HPC workflows, categorizing them into six execution motifs commonly used in scientific workflows: Steering, Multistage Pipeline, Inverse Design, Digital Replica, Distributed Models, and Adaptive Training. We highlighted prominent middleware exemplars of each motif and discussed their respective performance implications, revealing the performance challenges due to the complexity of integrating workflows across disparate resources. Moreover, we identify the main challenges of deploying AI-coupled HPC workflows at scale, such as the need for co-designed facilities and further development of middleware and benchmarks. We also identify opportunities for advancement in each area. Specifically, our survey identified a significant gap in the available workflow benchmarks to assess the performance of the six execution motifs. As such, we acknowledge one of the weaknesses in our survey is related to this dearth of benchmarks, which limited our performance discussions to simpler workflows. To address this issue, we propose a new suite of benchmarks designed to target each execution motif.
@@
AI-coupled HPC workflows also offer benefits for pre-training and fine-tuning of advanced AI systems. For example, in the context of the Adaptive Training motif, training LLMs at scale on HPC offers potentially more efficient alternatives at larger scales and lower costs than training in the cloud due to specific performance optimizations at scale. In general, the deployment of advanced AI systems within AI-coupled HPC workflows will benefit from the robust computational power of HPC systems.
@@
Workflows are becoming first-class applications for procuring future DOE HPC systems and implementing the Integrated Research Infrastructure (IRI) [73]. For example, the NERSC-10 RFP technical specifications [159] highlight three classes of workflows to support scientific research campaigns on the future HPC system: high-performance simulations and modeling workflows, high-performance AI workflows, and cross-facility workflows for real-time steering of experiments. These workflows will likely exhibit one or more or a combination of the AI-coupled HPC workflow motifs introduced in Section 2. Similarly, the technical specifications of OLCF-6 at the Oak Ridge National Laboratory [144] also emphasize the fact that scientific campaigns are inherently formulated as workflows that include “phases of HPC simulation, high-bandwidth communication, interleaved AI training with simulation and inferencing, and data processing”. In both cases, one of the main objectives in explicitly including workflows in the specifications is to move (part of) the management of such workflows from user space to the software ecosystem operated by the facilities. A closer and hopefully better integration with the platform should address the performance challenges related to AI-coupled HPC workflows that we listed in Section 4.
@@
The next step in scientific computing performance, scale, and sophistication will require AI-coupled HPC workflows to be a central component of future scientific campaigns. A thorough understanding of AI-coupled HPC workflows and the development of new benchmarks will create a strong foundation for addressing future advancements in the post-exascale era [161].

### v2

@@ 4.3.2 Multistage pipeline
In addition to dataflow across different “executables”, this gives rise to numerous performance issues. For example, the concurrent execution of multiple different stages with different computational costs and resource requirements results in a significant load balancing and workload management challenge. Furthermore, suppose AI models are used in conjunction with or as a substitute for each stage. In that case, many of the problems intrinsic to other motifs arise (e.g., dataflow performance bottlenecks in the Dynamic Orchestration motif).
@@ 4.3.3 Inverse design
Inverse design is a technique of calculating the causal factors that produced them from a set of observations. Such approaches in various domains include the search for catalysts [114], nanophotonics [115], and material design [116]. The performance challenges in inverse design stem from both the AI model’s capabilities and computational constraints, particularly when coupled with high-performance computing (HPC) systems. AI models must accurately represent causal factors from data, but their scalability and ability to handle high data rates, such as those from synchrotron facilities, present significant challenges. Computationally, traditional methods like pseudo-Voigt fitting are limited, and AI-based approaches like BraggNN [117] reduce bottlenecks, yet still require substantial resources. Coupling AI with HPC introduces additional issues, including efficient resource allocation across CPUs, GPUs, and accelerators, managing data transfer between nodes, and ensuring parallelism without performance degradation. Addressing these challenges is crucial to maintaining the efficiency and scalability of inverse design workflows in high-data environments.
@@ 4.3.4 Digital replica
The Digital Replica motif describes the integration of simulations with machine learning models, either in cognitive simulations (CogSim) or digital twins. The performance depends primarily on how the model is integrated into the simulation (coupling type) and compute heterogeneity. As mentioned earlier, this falls into two approaches: (1) in-memory or (2) distributed (via inference server). Jia et al. [118] used an in-memory inference approach via the TensorFlow C API to integrate a molecular dynamics (MD) simulation with a machine-learned model. Their model can efficiently scale up to the entire Summit supercomputer, achieving 91 PFLOPS (Peta FLoating-point Operations Per Second) in double precision and 162/275 PFLOPS in mixed-single/half-precision. For this reason, they won the Gordon Bell Prize in 2020.
@@
4.3.5 Distributed models
@@ 4.3.6 Adaptive execution for training large AI models
We discuss this motif for three types of implementations: Hyperparameter Optimization (HPO), Neural Architecture Search (NAS), and LLM training. HPO is typically used to optimize a set of training hyperparameters, such as batch size, learning rate, etc.; this is typically performed by training several neural networks in an asynchronous parallel fashion. Optimization methods generally fall into either random search or Bayesian camps, with novel augmentations such as Hyperband designed to speed up the evaluations by adaptive resource allocation and early stopping [121]. Tools such as KerasTuner may be run distributed using a chief-worker strategy to scale such methods across many nodes on HPC [122].
@@
LLMs are particularly challenging to train primarily because the models are generally much larger than what can fit within the memory of a single GPU, and the models need to be trained on colossal amounts of data. For example, GPT-3 has 175 billion trainable parameters and is trained on 499 billion tokens [128]; it has been estimated that it would take 288 years to train GPT-3 on a single V100 GPU [129]. Considerable developments in data and model parallelism have enabled significant speedup by using data-based parallelism methods, such as Horovod, and a combination of pipeline and tensor model parallelism [130]. Using such a triad strategy of data, tensor, and pipeline model parallelism along with DeepSpeed by Microsoft, Narayan et al. [129] have demonstrated the ability to train an LLM with one trillion parameters at 502 PF/s on 3,072 GPUs. Additionally, with the increasing number of non-traditional AI accelerators, training LLMs require different approaches to leverage the hardware. Emani et al. [131] provided a comparative evaluation of training LLM models on diverse AI accelerators. Recent studies have also further explored efficient training strategies and performance implications of up to a trillion-parameter LLMs on leadership-class supercomputers [132, 133]. Achieving such a feat required overcoming GPU memory limitations and minimizing inter-node communication latencies using a combination of pipeline and tensor model parallelism.
@@ 5 Open challenges and opportunities
The integration of AI into high-performance computing (HPC) workflows has unlocked new capabilities but also introduced complex challenges that span both computational and methodological dimensions. In this section, we identify four key challenges that must be addressed to advance AI-coupled HPC workflows. First, there is a critical need for co-designing hardware, middleware, and software to efficiently support both AI-coupled and traditional HPC workloads, ensuring optimal utilization of computing resources. Second, managing the massive and diverse datasets generated by AI-HPC workflows necessitates the development of scalable and efficient data handling strategies that address I/O bottlenecks, data movement, and real-time processing. Third, the absence of standardized benchmarking tools for evaluating AI-HPC workflows hinders systematic performance comparisons and optimization efforts across different execution motifs. Finally, enabling seamless interoperability, energy-efficient execution, and adaptive resource management remains a significant hurdle, requiring new approaches to dynamic workflow scheduling and cross-platform integration. Addressing these challenges presents opportunities to enhance the scalability, efficiency, and scientific impact of AI-HPC workflows, ultimately enabling more effective simulations, data analysis, and scientific discoveries.
@@ 5.2 Data and I/O Challenge
HPC platforms have settled on POSIX-compliant parallel file systems as the primary storage engine and interface. Even for an object-based solution such as DAOS, a POSIX shim layer is still needed to meet the requirement of more traditional modes in scientific applications. However, large, sequential read-and-write is no longer the dominant workload pattern for many data-intensive AI-coupled HPC workflows. Instead, small, random read/write is much more prevalent – which traditional POSIX-based parallel file systems have struggled to keep pace with. The nominal approach is to provision a node-local storage layer, where a small amount of SSD-backed devices are positioned close to the compute blades and generally outside the storage network used by the parallel file system. The challenge is that decoupling from primary storage necessitates data staging, which can be achieved through policy-driven data tiering or workflow support to streamline the process. The current solution space is not mature enough and thus renders this a particular bottleneck to overcome for data-intensive workflow integration. Projects such as NoPFS and SmartRedis discussed in Section 3 are attempts in the right direction. More recent OLCF-6 RFP [146] calls for an AI-optimized storage system and highlights the challenges and needs in this domain.
@@ 5.3 Benchmarking
To effectively evaluate the performance of AI-coupled HPC workflows, benchmarking them on diverse use case scenarios/motifs is essential. Such benchmarks could reveal potential performance bottlenecks and, hence, lead to optimized solutions. Several efforts in this direction include MLPerf [147], HPCAI500 [148], SAIBench from BenchCouncil [149], and HPL-AI [134]. MLPerf [147], developed by the MLCommons consortium of industry, academia, and research laboratories, is a state-of-the-art benchmark suite that includes various machine learning tasks. The MLPerf HPC benchmark [150] has been developed to represent scientific machine learning workloads typical of traditional HPC environments. There is also an ongoing effort to include AI-HPC workflow benchmarks in this suite in collaboration with the MLCommons Science group [151].
@@ 5.4 Cross-cutting: Interoperability, Energy-Efficiency and Resource Allocation
AI-coupled HPC workflows often require a high-level of interoperability, typically entailing connecting/interfacing with multiple endpoints for data, AI models, and workflow integration. There are two emerging patterns of choice: (1) REST API [153] and (2) message queue-based approach [82]. Each has pros and cons; currently, no widely accepted interfacing mechanisms exist. Further, time-sensitive patterns and long-term campaign patterns, as highlighted by the IRI blueprint report [71], demonstrate the need for interfacing beyond a single institution’s boundary, which brings a host of related challenges such as federated IDs, authentication, authorization, access control, etc., to implement a true end-to-end workflow integration. These challenges are more pronounced in motifs connecting experiments with HPC simulations (e.g., Distributed Models and Inverse Design) and in supporting Dynamic Orchestration and Multistage Pipeline motifs, which is only partially addressed by specialized frameworks such as MuMMI or DeepDriveMD, and more general frameworks such as SmartSim (ref. Section 3).
The importance of energy efficiency in high-performance computing (HPC) has intensified due to the hard limits on maximum power capacity. A critical observation in most AI-HPC workflows is their relatively low utilization of peak performance, which may be due to data movement or other I/O activities [96]. This sub-optimal utilization prompts a vital question: what should be the ultimate goal in this context? The answer might lie in optimizing how one could optimize productivity (Science/Watt) rather than raw performance (FLOP/Watt). This optimization raises a further dilemma: should the focus be on time-to-solution or throughput? One approach could be accelerating processes through high concurrency levels and GPU utilization. Alternatively, engaging GPUs less frequently might be more power-efficient, leading to a trade-off of lower throughput for energy conservation, albeit with longer processing times [154]. Other research areas for improving the energy efficiency of AI-coupled HPC workflows involve dynamic scheduling policies, data locality, ensemble pruning of unnecessary runs, and optimizations of convolution implementations [155]. Benchmarking such workflows will be key for understanding how to optimize energy efficiency.
@@ 6 Summary and Discussion
This paper surveyed the extensive landscape of AI-coupled HPC workflows, categorizing them into six execution motifs commonly used in scientific workflows: Dynamic Orchestration, Multistage Pipeline, Inverse Design, Digital Replica, Distributed Models, and Adaptive Training. We highlighted prominent middleware exemplars of each motif and discussed their respective performance implications, revealing the performance challenges due to the complexity of integrating workflows across disparate resources. Moreover, we identify the main challenges of deploying AI-coupled HPC workflows at scale, such as the need for co-designed facilities and further development of middleware and benchmarks. We also identify opportunities for advancement in each area. Specifically, our survey identified a significant gap in the available workflow benchmarks to assess the performance of the six execution motifs. As such, we acknowledge one of the weaknesses in our survey is related to this dearth of benchmarks, which limited our performance discussions to simpler workflows. To address this issue, we propose a new suite of benchmarks designed to target each execution motif.
@@
These integrated workflows also offer benefits for pre-training and fine-tuning of advanced AI systems. For example, in the context of the Adaptive Training motif, training LLMs at scale on HPC offers potentially more efficient alternatives at larger scales and lower costs than training in the cloud due to specific performance optimizations at scale. In general, the deployment of advanced AI systems within such workflows will benefit from the robust computational power of HPC systems.
@@
Workflows are becoming first-class applications for procuring future DOE HPC systems and implementing the Integrated Research Infrastructure (IRI) [71]. For example, the NERSC-10 RFP technical specifications [159] highlight three classes of workflows to support scientific research campaigns on the future HPC system: high-performance simulations and modeling workflows, high-performance AI workflows, and cross-facility workflows for real-time steering of experiments. These workflows will likely exhibit one or more or a combination of the AI-coupled HPC workflow motifs introduced in Section 2. Similarly, the technical specifications of OLCF-6 at the Oak Ridge National Laboratory [146] also emphasize the fact that scientific campaigns are inherently formulated as workflows that include “phases of HPC simulation, high-bandwidth communication, interleaved AI training with simulation and inferencing, and data processing”. In both cases, one of the main objectives in explicitly including workflows in the specifications is to move (part of) the management of such workflows from user space to the software ecosystem operated by the facilities. A closer and hopefully better integration with the platform should address the performance challenges related to AI-integrated workflows that we listed in Section 4.
@@
The next step in scientific computing performance, scale, and sophistication will require AI-coupled HPC workflows to be a central component of future scientific campaigns. A thorough understanding of these workflows and the development of new benchmarks will create a strong foundation for addressing future advancements in the post-exascale era [161].